OpenStudy (anonymous):

a/2+2/3=5/6 Find the LCD that would eliminate the fractions in this equation.

OpenStudy (solomonzelman):

Hint:\[2 \color{red} { \times 3 } =6\]\[3 \color{red} { \times 2 } =6\]\[A nd,~~~~~~~6 =6\]

OpenStudy (anonymous):

Find the LCD that would eliminate the fractions in this equation a/2+2/3=5/6

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf ■~Solution: }\) \(\LARGE\color{blue}{ \bf \frac{a}{2}+ \frac{2}{3}= \frac{5}{6} }\) \(\LARGE\color{blue}{ \bf \frac{a\color{red} { \times 3 } }{2\color{red} { \times 3 } }+ \frac{2\color{red} { \times 2 } }{3\color{red} { \times 2 } }= \frac{5}{6} }\) \(\LARGE\color{blue}{ \bf \frac{3a}{6}+ \frac{4}{6}= \frac{5}{6} }\) \(\LARGE\color{blue}{ \bf \frac{3a+4}{6}= \frac{5}{6} }\) \(\LARGE\color{blue}{ \bf 3a+4=5 }\) \(\LARGE\color{blue}{ \bf 3a=1 }\) \(\LARGE\color{blue}{ \bf a= \frac{1}{3}}\)